OpenStudy (anonymous):

The perimeter of a rectangle is 84 feet. the length of the garden is twice the width. What is the width of the garden?

OpenStudy (raden):

let : the its wide =w, so its length=2w

OpenStudy (raden):

width*

OpenStudy (anonymous):

What next

OpenStudy (raden):

P=2*l + 2*w P=perimeter

OpenStudy (raden):

change l=2w to that formula and put P=84

OpenStudy (anonymous):

Would the answer be 41?

OpenStudy (anonymous):

show your work @makenzieprater

OpenStudy (raden):

re-check again

OpenStudy (anonymous):

I dont understand how to do it

OpenStudy (anonymous):



OpenStudy (anonymous):

go all the way around, that's the perimeter... how many w's?

OpenStudy (anonymous):

2 w's

OpenStudy (anonymous):

all the way around is not 2w's

OpenStudy (anonymous):

go around and count them up.

OpenStudy (anonymous):

I mean 4

OpenStudy (anonymous):

keep guessing.

OpenStudy (anonymous):

or just count them.

OpenStudy (anonymous):

6?

OpenStudy (anonymous):

good guess it's 6 w's

OpenStudy (raden):

w+2w+2w+w= ... w?

OpenStudy (anonymous):

6w = 84

OpenStudy (anonymous):

@RadEn 6w

OpenStudy (anonymous):

so w= ..?

OpenStudy (anonymous):

6?

OpenStudy (anonymous):

because 6 times 6 is 84 right?

OpenStudy (anonymous):

Yes

OpenStudy (raden):

yes, solve for w like @Algebraic! said

OpenStudy (anonymous):

What od i do next

OpenStudy (anonymous):

if you work for 6 hours and make a total of 84 dollars, how much are you getting paid per hour?

OpenStudy (raden):

6w=84, just divide by 6 to both sides, what happend next ?

OpenStudy (anonymous):

14 an hour

OpenStudy (anonymous):

right.

OpenStudy (anonymous):

because 6*14 = 84

OpenStudy (anonymous):

so what is w?

OpenStudy (anonymous):

w=14

OpenStudy (anonymous):

yep.

OpenStudy (anonymous):

is that the answer then?

OpenStudy (anonymous):

you decide.

OpenStudy (raden):

look at ur question, to find its width or length ?